# FedTSA: A Cluster-based Two-Stage Aggregation Method for Model-heterogeneous Federated Learning

### v1

@@ Abstract
Despite extensive research into data heterogeneity in federated learning (FL), system heterogeneity remains a significant yet often overlooked challenge. Traditional FL approaches typically assume homogeneous hardware resources across FL clients, implying that clients can train a global model within a comparable time. However, in practical FL systems, clients often have heterogeneous resources, which impacts their capacity for training tasks. This discrepancy highlights the significance of exploring model-heterogeneous FL, a paradigm that allows clients to train different models based on their resource capabilities. To address this, we introduce FedTSA, a cluster-based two-stage aggregation method tailored for system heterogeneity in FL. FedTSA starts by clustering clients based on their capabilities, then conducts a two-stage aggregation, i.e., conventional weight averaging for homogeneous models as Stage 1, and deep mutual learning with a diffusion model for aggregating heterogeneous models as Stage 2. Extensive experiments not only show that FedTSA outperforms the baselines, but also explore various factors influencing model performance, thereby validating FedTSA as a promising approach for model-heterogeneous FL.
@@ 1 Introduction
Figure 1: Left: FL in traditional homogeneous settings. All clients share the same global model and the server performs weight averaging to aggregate them. Right: FL in heterogeneous settings. Due to the resource constraints, some clients can only train simple models, leading to heterogeneous models. As these models have different dimensions, how to aggregate them becomes the core challenge.
@@
A vanilla idea is to allow clients to adopt different model architectures tailored to their capabilities. For instance, a resource-constrained Raspberry Pi might use a two-layer convolutional neural network (CNN) model, while a GPU-equipped PC could employ a ResNet18 model [15]. When using traditional FL algorithms like FedProx [28], aggregating these diverse architectures becomes challenging due to the varying dimensions of model parameters, as illustrated in Figure 1. This discrepancy makes the weight averaging on the server infeasible. Knowledge distillation (KD)-based approaches have been proposed to mitigate this challenge by only aggregating and aligning with the global knowledge [24, 13]. However, they often require a public dataset on the client side, raising privacy and resource concerns [53]. Data-free KD-based methods have emerged as potential solutions to mitigate this limitation [58, 31], but they come with substantial communication costs and the issue of low-quality generated data. Another strategy is partial training, where each client trains a subnetwork of the global model [10, 1] and only the overlapping subnetworks will be aggregated. However, a critical question in partial training concerns the non-overlapping model parameters, which may contain important information but never get updated.
@@
In this paper, we propose FedTSA, a cluster-based two-stage aggregation method for model-heterogeneous FL. For any given FL training task, FedTSA starts with clustering clients according to their hardware resources using kernel density estimation (KDE) [21]. Afterward, FedTSA assigns varying pruning rates to different clusters, leading to heterogeneous models across clusters. Within each cluster, the homogeneous clients conduct local updates and perform model aggregation in conventional FL training styles, a process we refer to as Stage 1 aggregation. To aggregate heterogeneous models from Stage 1, we employ deep mutual learning (DML) [56] on the server based on the data generated by a diffusion model [17]. This phase is refer to as Stage 2 aggregation. The employment of the diffusion model not only obviates the need for crafting public datasets but also fosters the generation of superior-quality data, enhancing DML efficacy. Moreover, unlike FedGen [58], computation-intensive data generation tasks are offloaded to the server side, with clients only sending prompts to guide this process, which ensures a high degree of flexibility in training tasks.
Our extensive experiments on computer vision (CV) datasets demonstrate FedTSA’s effectiveness, outperforming other state-of-the-art model-heterogeneous FL approaches. We also provide an in-depth analysis of various factors influencing FedTSA, offering valuable insights for future model-heterogeneous FL research. Our contributions are threefold:
  • We introduce the novel FedTSA framework to enable model-heterogeneous FL. FedTSA first clusters clients according to their processing capabilities using KDE, then conducts a two-stage aggregation to aggregate heterogeneous models.
  • FedTSA innovatively leverages a diffusion model to generate data, thereby addressing the challenges of constructing a public dataset. This approach not only enhances data privacy but also significantly boosts the framework’s adaptability to various training tasks.
  • Extensive experimental evaluations on three CV datasets demonstrate FedTSA’s superior performance over existing baselines. Moreover, our investigations into influential factors present critical insights, further validating FedTSA’s effectiveness.
@@ 2 Related work
Model-heterogeneous FL. Different from the conventional FL paradigm [32, 49, 28, 48], model-heterogeneous FL [11] allows clients to train models of varying complexities based on their processing abilities, thus mitigating the system heterogeneity issues. Inspired by the idea of KD [16, 35], [24] proposes FedMD, a logit-based FL algorithm. Instead of transferring model parameters, FedMD employs logits as the communicative data medium and trains the global model by aggregating and approaching these logits. However, FedMD relies on a public dataset to align the input data, leading to privacy concerns and extra communication costs. This problem exists in several similar works, like [5, 20]. Subsequent works such as FedDF [31] and FedGen [58] address these limitations by either leveraging unlabeled data or training a generator to produce synthetic data to achieve data-free KD. However, FedDF requires prior knowledge to choose proper unlabeled data, while FedGen suffers from high computation cost as the generators are deployed and worked in each client. Moreover, whether it is unlabeled data or data generated by generative adversarial networks (GAN) [14], the quality of the data cannot be guaranteed, which will potentially lead to a decrease in model performance. Based on the idea of class prototype [41], FedProto [44] aggregates heterogeneous models by exchanging and aligning prototypes. In addition, HeteroFL [10] achieves model-heterogeneous FL by only aggregating the overlapping submodel parts, thus enabling clients to train submodels with different sizes. Despite existing methods adopt different strategies to achieve model-heterogeneous FL, none of them illustrate how to allocate different models to the clients.
Diffusion Models. With large-scale training data and model parameters, various text-to-image generative models are now capable of creating high-quality images [36, 9, 40]. Among them, diffusion models [17] have achieved significant success in the image synthesis domain and AI-generated content (AIGC) fields, such as the popular Stable Diffusion [38] and DALL-E2 [37]. A typical diffusion model includes a forward diffusion stage and a reverse diffusion stage [8]. Given a textual prompt, the diffusion model is able to convert Gaussian noise into a text-compliant image through an iterative denoising step. Despite diffusion models showcasing excellent performance, only a few works have leveraged them in FL. [51] and [57] combine diffusion models with FL to mitigate non-IID problems, while [45] applies FL to the training process to address privacy issues. FedTSA is the first work that leverages the advantages of diffusion models to achieve model-heterogeneous FL. Compared to other methods that either use a well-crafted public dataset or train a generator on the client side, employing a diffusion model not only improves the quality of the generated data but also offers greater flexibility in changing training tasks and better privacy protection.
@@ 3 Motivation Study
We motivate this study by presenting the impact of system heterogeneity through a small-scale experiment. We establish a heterogeneous system comprising our types of devices: three Raspberry Pi 4B units, three LattePanda 2 Alpha 864s [23], two PCs without GPU, and one PC equipped with an NVIDIA 3080Ti GPU. These devices range from embedded systems to PCs with GPUs, effectively presenting the system heterogeneity prevalent in real-world scenarios. The core specifications of each device type are detailed in Table 2. Given the limited resources of the Raspberry Pi, we select a FL training task using the MNIST dataset and a CNN model. This model includes two convolutional layers with max pooling, followed by two fully connected layers. Each device serves as a FL client, with an additional machine serving as the central server for aggregation. We employ FedAvg [32] for this study.
@@
We design four experiments to evaluate the impact of system heterogeneity. The settings and the wall-clock convergence times are detailed in Table 2. We can observe that when all devices participate in training, the convergence speed is very slow, exceeding 2700 seconds. This is attributed to tail latency [25] caused by the existence of Raspberry Pi devices, which delay the process even after other devices have completed their training. The convergence times for the remaining three experiments are reduced to 1428, 210, and 143 seconds, respectively, showing decreases of 47%, 92% and 95% compared to the initial scenario. These results show the idle times of faster devices, highlighting the significant impact of system heterogeneity on FL training efficiency.
@@ 4.1 Problem Definition
In this work, we consider a system-heterogeneity setting in FL. Suppose we have $n$ clients with their private local dataset $D_{i}$ from distribution $P_{i}(x,y)$, with $x$ and $y$ representing the input data and their respective labels. Unlike most works that assume clients have similar hardware resources, we consider the presence of heterogeneous resource distributions among clients, denoted as $R_{1},R_{2},...,R_{n}$. Due to this challenge, a standardized global model $W_{g}$ loses its applicability, as clients with limited resources cannot afford the computational costs if $W_{g}$ is complex. Therefore, the problem becomes how to enable clients to deploy different models $W_{1},W_{2},...,W_{n}$ according to their respective resources $R_{1},R_{2},...,R_{n}$, while still achieve FL training on the server side. Similar to conventional FL approaches, for the $i^{th}$ client, the training procedure is to minimize the loss as follows:
@@ 4.2 Two-Stage Aggregation Framework
Stage 2 in Figure 2 depicts how models with different architectures achieve aggregation with the help of DML and a diffusion model. Here, data generated from the diffusion model enables knowledge transfer between heterogeneous models. Specifically, each model initially takes these data as inputs to generate logits, which represent the output vectors of the neural network before applying the softmax function. Afterward, the models perform gradient descent using the same generated data with Kullback Leibler (KL) divergence loss. Contrary to conventional KD-based FL methods, in our setting, no global model exists on the server side. Therefore, KL divergence loss is calculated by comparing each set of logits with a consensus. In this context, the consensus is defined as the average of all logits generated by different models. After the DML process, knowledge is interchanged among models, finishing Stage 2 aggregation. The updated models are sent back to their respective clusters for further updates. Algorithm 1 outlines the FedTSA learning process, where the LocalUpdate function is the same one as FedAvg.
@@ 5.1 Experimental Setup
Datasets and Models. We evaluate the performance of FedTSA using three benchmark datasets, i.e., CIFAR-10, CIFAR-100 [22], and Tiny-ImageNet [7]. Different data distributions are considered in the experiments. In the IID setting, data is evenly distributed to each client in terms of both quantity and class distribution. In the non-IID settings, we employ the Dirichlet distribution $Dir(\alpha)$ [26] to allocate the dataset among the clients, where a smaller $\alpha$ indicates higher data heterogeneity. We set the $\alpha$ to 0.3 and 0.6 to introduce different data heterogeneity levels. We use the ResNet18 as the model in the experiments. To incorporate model heterogeneity, we adjust the model complexity by varying the hidden sizes of the residual blocks with a pruning rate, where a smaller rate represents a simpler model.
@@ 5.2 Performance Evaluation and Analysis
Performance Comparison. From the results presented in Table 3, we can observe that FedTSA outperforms other FL baselines across both IID and non-IID settings, highlighting its promising performance and robustness to data heterogeneity. An exception is noted in the IID setting of CIFAR-10 dataset, where FedTSA slightly lags behind the upper bound performance of FedAvg. However, we should notice that the upper bounds for FedAvg and FedProx are based on an idealized scenario where all clients have unpruned models. Given this context, FedTSA still maintains a dominant performance over most upper bounds across both IID and non-IID settings. This superior performance is attributed to FedTSA’s innovative strategy. Contrasting with algorithms that conventionally replace the global model with a local model for round updates, FedTSA fosters inter-client learning through DML. Specifically, FedTSA utilizes the KL divergence loss to learn the average logits from other client models, thereby effectively obviating the need to rely on a singular global model. Consequently, each model not only gains knowledge from its counterparts but also sidesteps potential information loss that may arise from direct replacement. Besides, we can notice that FedProto presents the poorest accuracy, revealing that prototype-based FL methods do not perform as well as methods based on knowledge distillation and local training in the presence of model heterogeneity. The underlying reason is that simpler models produce prototypes of lower quality, which in turn degrades the quality of the aggregated global prototypes, ultimately leading to poor model training outcomes.
@@
Effect of Different Degrees of Model Heterogeneity. We evaluate the performance of FedTSA across varying degrees of model heterogeneity, comparing it with other model-heterogeneous FL baselines. To introduce model heterogeneity, we continue to utilize the pruning rate and keep the number of clients to 20. We select CIFAR-10 as the dataset and configure the data distribution to be IID. The specific settings for heterogeneity and the corresponding experimental results are detailed in Table 4.
@@
We can observe that our method FedTSA outperforms the baseline methods across all five degrees of heterogeneity. This performance illustrates FedTSA’s robust capability in knowledge exchange within heterogeneous model environments. Notably, the accuracy of HeteroFL decreases significantly when the heterogeneity degrees shift from [0.6, 0.8, 1.0] to [0.1, 0.4, 0.7, 1.0]. The decline can be attributed to HeteroFL’s reliance on a partial training mechanism, where only the common parts of heterogeneous models are aggregated. Therefore, if the pruning rate is excessively low, e.g., 0.1, the resulting model becomes overly simplistic, failing to effectively learn data patterns and thus contributing minimally to the global model. In contrast, FedTSA gains contributions from heterogeneous models through DML, where models with larger pruning rates can guide those with smaller ones, just like KD. That is the reason why FedTSA does not suffer from a significant decrease in accuracy. By effectively harnessing the strengths of models across a spectrum of heterogeneity, FedTSA demonstrates superior adaptability and efficiency in learning from diverse model architectures, thereby enhancing overall performance in model-heterogeneous FL scenarios.
@@
6 Discussion
FedTSA enables clients to send prompts that help generate data for the diffusion model. Concerns may arise regarding privacy, particularly the possibility of servers knowing the details of training tasks from these prompts. In fact, it is common practice for servers to be cognizant of the tasks being executed by clients. This knowledge is essential for the server to select the appropriate model architecture and initialize the model. Despite being aware of the task type, the server does not need to know the specific data of the clients, still achieving privacy protection. Furthermore, it is not mandatory for the label information within a prompt to match an actual label from the local dataset. Any label related to the training task can be used. The primary objective of the data generation process is to align the logits produced by heterogeneous models, facilitating their aggregation. This approach also addresses concerns regarding discrepancies between generated data and the actual data of certain clients. The generated data only serves as a bridge in Stage 2, ensuring the transfer of knowledge across heterogeneous models.
Despite FedTSA showing excellent performance, some trade-offs exist. First, the diffusion model introduces additional computational overhead to generate image data. However, it also offers better flexibility to switch different training tasks by simply adjusting the prompts than fixed public datasets. Second, like other existing model-heterogeneous FL methods, FedTSA has been only verified on CV tasks. The adaption and experiments of FedTSA on natural language processing tasks are our main effort direction in the future.
7 Conclusion
This paper proposes FedTSA, a novel cluster-based two-stage aggregation method for model-heterogeneous FL. FedTSA first conducts clustering to generate heterogeneous cluster-based models, then leverages DML on the server side to aggregate these models based on the data generated by a diffusion model. Experimental results on CV datasets show that FedTSA surpasses baselines in both IID and non-IID settings. Further experiments on the influential factors evaluate the effect of these factors on model training. In future research, we will investigate methods for achieving model-heterogeneous FL without any external data, and make FedTSA available for more scenarios, such as natural language processing and multimodal tasks.

### v2

@@ Abstract
Despite extensive research into data heterogeneity in federated learning (FL), system heterogeneity remains a significant yet often overlooked challenge. Traditional FL approaches typically assume homogeneous hardware resources across FL clients, implying that clients can train a global model within a comparable time frame. However, in practical FL systems, clients often have heterogeneous resources, which impacts their training capacity. This discrepancy underscores the importance of exploring model-heterogeneous FL, a paradigm allowing clients to train different models based on their resource capabilities. To address this challenge, we introduce FedTSA, a cluster-based two-stage aggregation method tailored for system heterogeneity in FL. FedTSA begins by clustering clients based on their capabilities, then performs a two-stage aggregation: conventional weight averaging for homogeneous models in Stage 1, and deep mutual learning with a diffusion model for aggregating heterogeneous models in Stage 2. Extensive experiments demonstrate that FedTSA not only outperforms the baselines but also explores various factors influencing model performance, validating FedTSA as a promising approach for model-heterogeneous FL.
@@ 1 Introduction
Figure 1: Left: FL in homogeneous settings. All clients share the same global model and the server performs weight averaging to aggregate them. Right: FL in heterogeneous settings. Due to the resource constraints, some clients can only train simple models, leading to heterogeneous models. As these models have different dimensions, how to aggregate them becomes the core challenge.
@@
A vanilla idea is to allow clients to adopt different model architectures tailored to their capabilities. For instance, a resource-constrained Raspberry Pi might use a two-layer convolutional neural network (CNN) model, while a GPU-equipped PC could employ a ResNet18 model [15]. When using traditional FL algorithms, such as FedProx [28], aggregating these diverse architectures becomes challenging due to the varying dimensions of model parameters, as illustrated in Figure 1. This discrepancy makes the weight averaging on the server infeasible. Knowledge distillation (KD)-based approaches have been proposed to mitigate this challenge by only aggregating and aligning with the global knowledge [24, 13]. However, they often require a public dataset on the client side, raising privacy and resource concerns [53]. Data-free KD-based methods have emerged as potential solutions to mitigate this limitation [58, 31], but they come with substantial communication costs and the issue of low-quality generated data. Another strategy is partial training, where each client trains a subnetwork of the global model [10, 1] and only the overlapping subnetworks will be aggregated. However, a critical issue in partial training is the non-overlapping model parameters, which may contain important information but never get updated.
@@
In this paper, we propose FedTSA, a cluster-based two-stage aggregation method for model-heterogeneous FL. For any given FL training task, FedTSA starts with clustering clients according to their hardware resources using kernel density estimation (KDE) [21]. Afterwards, FedTSA assigns varying pruning rates to different clusters, leading to heterogeneous models across clusters. Within each cluster, the homogeneous clients conduct local updates and perform model aggregation in conventional FL training styles, a process we refer to as Stage 1 aggregation. To aggregate heterogeneous models from Stage 1, we employ deep mutual learning (DML) [56] on the server based on the data generated by a diffusion model [17]. This phase is refer to as Stage 2 aggregation. The employment of the diffusion model not only obviates the need for crafting public datasets but also fosters the generation of superior-quality data, enhancing DML efficacy. Moreover, unlike FedGen [58], computation-intensive data generation tasks are offloaded to the server side, with clients only sending prompts to guide this process, which ensures a high degree of flexibility in training tasks.
Our extensive experiments on computer vision (CV) datasets demonstrate FedTSA’s effectiveness, outperforming other state-of-the-art model-heterogeneous FL approaches. We also provide an in-depth analysis of various factors influencing FedTSA, offering valuable insights for future model-heterogeneous FL research. Our contributions are threefold: 1) We introduce the novel FedTSA framework to enable model-heterogeneous FL. FedTSA first clusters clients according to their processing capabilities using KDE, then conducts a two-stage aggregation to aggregate heterogeneous models. 2) FedTSA innovatively leverages a diffusion model to generate data, thereby addressing the challenges of constructing a public dataset. This approach not only enhances data privacy but also significantly boosts the framework’s adaptability to various training tasks. 3) Extensive experimental evaluations on three CV datasets demonstrate FedTSA’s superior performance over existing baselines. Moreover, our investigations into influential factors present critical insights, further demonstrating FedTSA’s effectiveness.
@@ 2 Related work
Model-heterogeneous FL. Different from the conventional FL paradigm [32, 49, 28, 48], model-heterogeneous FL [11] allows clients to train models of varying complexities based on their processing abilities, thus mitigating the system heterogeneity issue. Inspired by the idea of KD [16, 35], [24] proposes FedMD, a logit-based FL algorithm. Instead of transferring model parameters, FedMD employs logits as the communicative data medium and trains the global model by aggregating and approaching these logits. However, FedMD relies on a public dataset to align the input data, leading to privacy concerns and extra communication costs. This problem exists in several similar works [5, 20]. Subsequent works such as FedDF [31] and FedGen [58] address these limitations by either leveraging unlabeled data or training a generator to produce synthetic data to achieve data-free KD. However, FedDF requires prior knowledge to choose proper unlabeled data, while FedGen suffers from high computation cost as the generators are deployed and worked in each client. Moreover, the quality of unlabeled data or data generated by generative adversarial networks (GAN) [14] cannot be guaranteed, potentially leading to a decrease in model performance. Based on the idea of class prototype [41], FedProto [44] aggregates heterogeneous models by exchanging and aligning prototypes. In addition, HeteroFL [10] achieves model-heterogeneous FL by only aggregating the overlapping submodel parts, thus enabling clients to train submodels with different sizes. Despite existing methods adopt different strategies to achieve model-heterogeneous FL, none of them illustrate how to allocate different models to the clients.
Diffusion Models. With large-scale training data and model parameters, various text-to-image generative models are now capable of creating high-quality images [36, 9, 40]. Among them, diffusion models [17] have achieved significant success in the image synthesis domain and AI-generated content (AIGC) fields, such as the popular Stable Diffusion [38] and DALL-E2 [37]. A typical diffusion model includes a forward diffusion stage and a reverse diffusion stage [8]. Given a textual prompt, the diffusion model is able to convert Gaussian noise into a text-compliant image through an iterative denoising step. Despite diffusion models showcasing excellent performance, only a few works have leveraged them in FL. [51] and [57] combine diffusion models with FL to mitigate non-IID problems, while [45] applies FL to the training process to address privacy issues.
FedTSA is the first work that leverages the advantages of diffusion models to achieve model-heterogeneous FL. Compared to other methods that either use a well-crafted public dataset or train a generator on the client side, employing a diffusion model not only improves the quality of the generated data but also offers greater flexibility in changing training tasks and better privacy protection.
@@ 3 Motivation Study
We motivate this study by presenting the impact of system heterogeneity through a small-scale experiment. We establish a heterogeneous system comprising four types of devices, including three Raspberry Pi 4B units, three LattePanda 2 Alpha 864s [23], two PCs without GPU, and one PC equipped with an NVIDIA 3080Ti GPU. These devices range from embedded systems to PCs with GPUs, effectively presenting the system heterogeneity prevalent in real-world scenarios. The core specifications of each device type are detailed in Table 2. Given the limited resources of the Raspberry Pi, we select a FL training task using the MNIST dataset and a CNN model. This model includes two convolutional layers with max pooling, followed by two fully connected layers. Each device serves as a FL client, with an additional machine serving as the central server for aggregation. We employ FedAvg [32] in this study.
@@
We design four experiments to assess the impact of system heterogeneity. The settings and the wall-clock convergence times are detailed in Table 2. We can observe that when all devices participate in training, the convergence speed is extremely slow, exceeding 2700 seconds. This is attributed to tail latency [25] caused by the existence of Raspberry Pi devices, which delay the process even after other devices have completed their training. The convergence times for the remaining three experiments are reduced to 1428, 210, and 143 seconds, respectively, showing decreases of 47%, 92% and 95% compared to the initial scenario. These results show the idle times of faster devices, highlighting the significant impact of system heterogeneity on FL training efficiency.
@@ 4.1 Problem Definition
In this work, we consider a system heterogeneity setting in FL. Suppose we have $n$ clients with their private local dataset $D_{i}$ from distribution $P_{i}(x,y)$, with $x$ and $y$ representing the input data and their respective labels. Unlike most works that assume clients have similar hardware resources, we consider the presence of heterogeneous resource distributions among clients, denoted as $R_{1},R_{2},...,R_{n}$. Due to the challenge of heterogeneity, a standardized global model $W_{g}$ loses its applicability, as clients with limited resources cannot afford the computational costs if $W_{g}$ is complex. Therefore, the problem becomes how to enable clients to deploy different models $W_{1},W_{2},...,W_{n}$ according to their respective resources $R_{1},R_{2},...,R_{n}$, while still achieve FL training on the server side. Similar to conventional FL approaches, for the $i^{th}$ client, the training procedure is to minimize the loss as follows:
@@ 4.2 Two-Stage Aggregation Framework
Stage 2 in Figure 2 depicts how models with different architectures achieve aggregation with the help of DML and a diffusion model. Here, data generated from the diffusion model enables knowledge transfer between heterogeneous models. Specifically, each model initially takes these data as inputs to generate logits, which represent the output vectors of the neural network before applying the softmax function. Afterwards, the models perform gradient descent using the same generated data with Kullback Leibler (KL) divergence loss. Contrary to conventional KD-based FL methods, in our setting, no global model exists on the server side. Therefore, KL divergence loss is calculated by comparing each set of logits with a consensus. In this context, the consensus is defined as the average of all logits generated by different models. After the DML process, knowledge is interchanged among models, finishing Stage 2 aggregation. The updated models are sent back to their respective clusters for further updates. Algorithm 1 outlines the FedTSA learning process, where the LocalUpdate function is the same as FedAvg.
@@ 5.1 Experimental Setup
Datasets and Models. We evaluate the performance of FedTSA using three benchmark datasets, i.e., CIFAR-10, CIFAR-100 [22], and Tiny-ImageNet [7]. We consider both IID and non-IID data distributions in the experiments. In the IID setting, data is evenly distributed to each client in terms of both quantity and class distribution. In the non-IID settings, we employ the Dirichlet distribution $Dir(\alpha)$ [26] to allocate the dataset among the clients, where a smaller $\alpha$ indicates higher data heterogeneity. We set the $\alpha$ to 0.3 and 0.6 to introduce different data heterogeneity levels. We use the ResNet18 as the model in the experiments. To incorporate model heterogeneity, we adjust the model complexity by varying the hidden sizes of the residual blocks with a pruning rate, where a smaller rate represents a simpler model.
@@ 5.2 Performance Evaluation and Analysis
Performance Comparison. From the results presented in Table 3, we can observe that FedTSA outperforms other FL baselines across both IID and non-IID settings, highlighting its promising performance and robustness to heterogeneous data. An exception is noted in the IID setting of CIFAR-10 dataset, where FedTSA slightly lags behind the upper bound performance of FedAvg. However, we should notice that the upper bounds for FedAvg and FedProx are based on an idealized scenario where all clients have unpruned models. Given this context, FedTSA still maintains a dominant performance over most upper bounds across both IID and non-IID settings. This superior performance is attributed to FedTSA’s innovative strategy. Contrasting with algorithms that conventionally replace the global model with a local model for round updates, FedTSA fosters inter-client learning through DML. Specifically, FedTSA utilizes the KL divergence loss to learn the average logits from other client models, thereby effectively obviating the need to rely on a singular global model. Consequently, each model not only gains knowledge from its counterparts but also sidesteps potential information loss that may arise from direct replacement. Besides, we can notice that FedProto presents the lowest accuracy, revealing that prototype-based FL methods do not perform as well as methods based on knowledge distillation and local training in the presence of model heterogeneity. This is because simpler models produce lower-quality prototypes, which in turn degrades the quality of the aggregated global prototypes, ultimately leading to poor model training outcomes.
@@
Effect of Different Degrees of Model Heterogeneity. We evaluate the performance of FedTSA across varying degrees of model heterogeneity, comparing it with other model-heterogeneous FL baselines. To introduce model heterogeneity, we continue to use the pruning rate and keep the number of clients to 20. We select CIFAR-10 as the dataset and configure the data distribution to be IID. The specific settings for heterogeneity and the corresponding experimental results are detailed in Table 4.
@@
We can observe that FedTSA outperforms the baseline methods across all five degrees of heterogeneity. This performance illustrates FedTSA’s robust capability in knowledge exchange within heterogeneous model environments. Notably, the accuracy of HeteroFL decreases significantly when the heterogeneity degrees shift from [0.6, 0.8, 1.0] to [0.1, 0.4, 0.7, 1.0]. The decline can be attributed to HeteroFL’s reliance on a partial training mechanism, where only the common parts of heterogeneous models are aggregated. Therefore, if the pruning rate is excessively low, e.g., 0.1, the resulting model becomes overly simplistic, failing to effectively learn data patterns and thus contributing minimally to the global model. In contrast, FedTSA gains contributions from heterogeneous models through DML, where models with larger pruning rates can guide those with smaller ones, just like KD. That is the reason why FedTSA does not suffer from a significant decrease in accuracy. By effectively harnessing the strengths of models across a spectrum of heterogeneity, FedTSA demonstrates superior adaptability and efficiency in learning from diverse model architectures, thereby enhancing overall performance in model-heterogeneous FL scenarios.
@@
6 Conclusion and Discussion
This paper contributes FedTSA, a novel cluster-based two-stage aggregation method for model-heterogeneous FL. FedTSA first conducts clustering to generate heterogeneous cluster-based models, then leverages DML on the server side to aggregate these models based on the data generated by a diffusion model. This approach allows each client to train a model that aligns with its available resources, thereby avoiding straggler issues and enhancing overall efficiency. Experimental results on CV datasets show that FedTSA surpasses baselines in both IID and non-IID settings. Further experiments analyze the impact of various influential factors on model training, providing a deeper understanding of their effects.
Despite the excellent performance of FedTSA, some trade-offs exist. First, the diffusion model introduces additional computational overhead to generate image data. However, it offers better flexibility in switching training tasks by simply adjusting the prompts compared to fixed public datasets. We will explore methods for achieving model-heterogeneous FL without relying on external data. Second, like other existing model-heterogeneous FL methods, FedTSA has only been verified on CV tasks. Adapting and experimenting with FedTSA on natural language processing and multimodal tasks is a primary focus for future work.